Molecular weight of NaCl determined by freezing point depression of its 0.5 %aqueous solution is 30. The Apparent degree of dissociation of NaCl is..... 
Ans is 0.95... Plz explain the solution 

Grade:12

1 Answers

Umakant biswal
7 years ago
@ nik 
can u able to recall a formula in the solution chapter , the formula is 
delta Tf = i*Kf*m.
where t(f) is the freezing point , i is the vant hoff factor kf is the depression constant and m is the molecular mass, 
this formula  will give iinformation about dissociation.
and i = normal molar mass/observed molar mass = 58.5/30 = 1.95 = 1+ alpha
so, by applying the value and calculating we have 
alpha = 0.95 
HOPE IT CLEARS YOUR DOUBT 

ALL THE BEST ..